Stability investigation by means of semi-definite Lyapunov functions

Alexander Ignatyev Institute of Applied Mathematics and Mechanics, Ukraine

A nonautonomous system of ordinary differential equations is considered. This system has the zero solution, and there exists a nonnegative Lyapunov function which derivative is nonpositive. Theorems on the uniform stability and asymptotic stability are proved.